Husky and Starch Books provided the following information: Average selling price per book $25 Average variable...

  1. Husky and Starch Books provided the following information:

Average selling price per book $25

Average variable cost per book $13

Monthly fixed costs $6,000

Compute the following:

What is Husky and Starch’s contribution margin per unit in dollars?

What is Husky and Starch’s contribution margin ratio?

How many books must Husky and Starch sell to breakeven?

How many books must Husky and Starch sell to earn $9,000 per month?
Answer #1

1. Contribution margin per unit = Selling price - Variable cost = $25 - $13 = $12

2. Contribution margin ratio = ( Contribution margin per unit / Selling price) * 100 = (12 / 25) * 100 = 48%

3. Break even books = Fixed cost / Contribution margin per unit = 6,000 / 12 = 500 books

4. Books to earn $9,000 profit = (Fixed cost + Desired profit) / Contribution margin per unit

= (6000 + 9000) / 12 = 15,000 / 12 = 1250 books